### v2

@@ Abstract
Aims. Our goal is to study the gravitational effects caused by the passage of the Large Magellanic Cloud (LMC) in its orbit on the stellar halo of the Milky Way (MW).
Methods. We employed the Gaia Data Release 3 to construct a halo tracers data set consisting of K-Giant stars and RR-Lyrae variables. Additionally, we have compared the data with a theoretical model to estimate the DM subhalo mass.
Results. We have improved the characterisation of the local wake and the collective response due to the LMC orbit. On the other hand, we have estimated for the first time the dark subhalo mass of the Large Magellanic Cloud, of the order of $1.7\times 10^{11}$ M, comparable to previously reported values in the literature.
@@ 1 Introduction
Dark matter (DM) is central to the standard cosmological model (LCDM), providing gravitational support for forming the galaxies and systems of galaxies (Mo et al., 2010). Its existence is backed by a plethora of observational data, including galaxies’ rotation curves (Zwicky, 1933; Rubin & Ford, 1970), strong and weak gravitational lensing effects (Massey et al., 2010; Clowe et al., 2006), and even the presence of the baryonic acoustic oscillation (Planck Collaboration et al., 2020) in the earlier gravitational wells observed in the cosmic microwave background (CMB).
@@
Despite of these successes, we still lack of precise detection in the laboratory (Bernabei et al., 2022; XENON Collaboration et al., 2023; Barberio et al., 2022; Amaré et al., 2022) or by indirect astrophysical observation (Abdalla et al., 2022; Acharyya et al., 2023; Abe et al., 2023). Many theoretical candidates arise from physics beyond the standard model (BSMP), like Weakly Interacting Massive Particles (WIMPS). These massive ($m_{DM}\sim 100$ GeV) particles could weakly interact with nucleons, and therefore signals are looked at by several laboratory and accelerator experiments. Also, its annihilation signals could be detectable through $\gamma$-ray emission by high-energy telescopes. Despite a massive experimental effort, DM remains a theoretical hypothesis, albeit one with impressive empirical support.
@@
Other DM candidates could be more massive, such as primordial black holes (PBHs) (Carr & Kühnel, 2020) which were recently constrained with a series of consistency tests. Nowadays, there is still room to be an essential contributor to the DM content but limited to small windows in mass (Villanueva-Domingo et al., 2021). Other candidates include massive ultralight particles (ULDM) that could reach masses as low as $10^{-23}$ eV (Hui et al., 2017). Therefore the possible DM mass range remains unconstrained today. Additionally, DM particles could interact with themselves, have or have no spin, and other properties, including their mass, remain elusive.
@@
Depending on the nature of the dark matter particle, there are relevant changes in the structure and number of dark matter haloes and subhalos (Zavala & Frenk, 2019). For example, some candidates, like warm dark matter (WDM), introduce a cut-off scale in the initial power spectrum of mass fluctuations ($m_{DM}\sim 1$ KeV), and others a scale during the non-linear evolution phase where the DM particles self-interact (SIDM) (Tulin & Yu, 2018). Both processes change the abundance of dark matter subhalos and the density profiles of the DM halos in comparison to the predictions of the Cold Dark Matter (CDM) model.
@@
Buschmann et al. (2018) showed, using CDM simulations, that the gravitational pull of dark matter subhalos affects the distribution of stars in galactic haloes, and that could be used to discover dark subhalos (those without star formation in situ, a precise prediction of the CDM model) and also allows to test the nature of the DM particles itself. This work simulated a passing dark matter subhalo’s perturbation to the phase space stellar distribution. Stars are pulled towards the subhalo as it passes, leaving a distinctive feature in halo stars’ velocity and number density, known as a wake. This phenomenon was previously analytically described by Weinberg (1986) due to the gravitational friction that provokes the orbital decay of the satellite galaxies that inhabit the DM subhalo. There are some efforts to quantify the magnitude of phase-space perturbations caused by the passage of dark matter subhalos using simulations and their possible detection (Bazarov et al., 2022).
@@
Garavito-Camargo et al. (2019) using CDM simulations quantified the impact of the LMC’s passage on the density and kinematics of the Milky Way’s (MW) DM halo and the observability of these structures in the MW’s stellar halo. Their results indicated a pronounced wake, which could be decomposed in a Transient and a Collective response in both the DM and stellar halos distributions. Such effect was observed for the first time in the Milky Way halo stars by Conroy et al. (2021). The authors studied the effects induced by the Magellanic Clouds system merging on selected samples of Milky Way stars with precise Gaia Satellite measurements. This detection, and the increasing availability of stellar data from Gaia DR3 release, pave the way to a more precise measurement of the effect. Measurements of the wake on the perturber systems of reference will allow pursuit testing the DM particle nature using this merger data, as was proposed recently by Foote et al. (2023) (in the context of systems of galaxies see Furlanetto & Loeb (2002); Buehler & Desjacques (2023)). Furthermore, Aguilar-Argüello et al. (2022) and Cunningham et al. (2020) conducted studies on the decomposition into spherical harmonics of both density and velocity, respectively, in order to quantify the response of the dark matter halo to the passage of the Large Magellanic Cloud.
@@
Based on the findings of Conroy et al. (2021), we have used the Gaia Data Release 3 to study the DM subhalo of the Magellanic Clouds and our code developments will allow us to apply the methodology to other MW satellites galaxies and Globular clusters and even to develop methods to detect the presence of the dark subhalos predicted by the CDM model.
@@
This work is organised as follows. In Section 2, we present the data samples selection methodology used to identify the effects of the DM subhalo of the LMC on the MW stellar halo. In Section 3, we briefly describe the theoretical model. Meanwhile, we present our results in Section 4. Section 5 presents the conclusions and future perspectives. Finally, in Appendix A we present a description of the coordinates transformation, in Appendix B the machine learning algorithm used to estimate radial velocities and in Appendix C validation of our method to infer distances.
@@ 2 Data reduction
We studied the gravitational response of the Milky Way’s Halo to the passage of the Large Magellanic Cloud in its orbit. To achieve this, we used the data from the Gaia Data Release 3 (Gaia Collaboration et al., 2022, 2016) and created two catalogues of Halo tracers, namely the K-Giants and RR-Lyraes stars. We followed the steps proposed by Conroy et al. (2021) to address this task.
@@ 2.1 K-Giant data set
To construct the K-Giant catalogue, we start the analysis with 162240774 sources characterised by $ruwe$ values below 1.4, parallax measurements lower than 0.2 mas and galactic latitude $|b|>10^{\circ}$ to remove the galactic plane. To ensure data quality, we performed a series of cleaning procedures. First, we eliminated sources lacking of proper motion and photometric data. To account for dust extinction, we obtained the dustmap from Green (2018) and considered the SFD map to derive the excess colour, $E(B-V)$. Subsequently, we discarded all sources with $E(B-V)>0.3$. To obtain the corrected magnitudes, we considered the following coefficients: $A_{G}/E(B-V)=2.4$, $A_{BP}/E(B-V)=2.58$, and $A_{RP}/E(B-V)=1.65$. To focus solely on the giant branch, we restricted the selection to sources satisfying the condition $21.4<(BP^{*}-RP^{*})<$ 2, where $BP^{*}$ and $RP^{*}$ represent the corrected magnitudes. Next, following Riello et al. (2021), we performed the $3\sigma$ cut upon the corrected BP and RP flux excess factor ($C^{*}$). After completing the data-cleaning process, to ensure the purity of our catalogue specifically for K-Giant stars, we performed a cross-match with the spectral types provided by Gaia (gaiadr3.astrophysical parameters), we obtained a data set of 490669 sources. Finally, we restricted our analysis to objects within a galactocentric distance between $30$ kpc and $100$ kpc, remaining 245086 sources. Among them, only 10989 had measured radial velocity from Gaia (Katz et al., 2023).
@@
To validate our distance inference method we compared our calculated distances with reference values for some globular clusters, the Magellanic Clouds (LMC and SMC) and some satellite galaxies (see Appendix C for details). Our method successfully reproduced tabulated distances, since the mean calculated distances differed by less than $20\%$.
Afterwards, we implemented several masks to exclude known objects from our analysis. Specifically, we applied angular and distances masks for all the globular clusters listed in Harris (1996a) and all the satellite galaxies reported in Drlica-Wagner et al. (2020a).
Following the methodology proposed by Conroy et al. (2021), we employed proper motions to eliminate structures linked to the Sagittarius stream. To achieve this, we initially correct the proper motions due to the solar reflex motion, with the ”gala” package (Price-Whelan, 2017; Price-Whelan et al., 2020). The used parameters were $R_{\odot}=8.122$ kpc (GRAVITY Collaboration et al., 2019), $(V_{R,\odot},V_{\phi,\odot},V_{Z,\odot})=(-12.9,245.6,7.78)$ km/s (Drimmel & Poggio, 2018) and the distance of the Sun from the Galactic mid-plane $Z_{\odot}=20.8$ pc (Bennett & Bovy, 2019). For $b>10^{\circ}$ and $|B_{Sgr}|<15^{\circ}$, where $B_{Sgr}$ is the latitude in the frame of Sagittarius orbital plane, we remove part of the northern arm of the stream by taking out the sources that have $\mu_{\alpha*}>-1.3$ mas/yr, $-0.4<\mu_{\delta}<0.3$ mas/yr and $\mu_{\delta}>1.7\mu_{\alpha*}+0.4$. To eliminate the rest of the north arm, we applied a mask to the region with coordinates $b>0^{\circ}$ and $180^{\circ}<l<210^{\circ}$. The final selection of sources is based on proper motions, that is keeping only those that satisfy $\mu_{\alpha*}^{2}+(\mu_{\delta}+0.1)^{2}<0.5^{2}$ (Conroy et al., 2021). By implementing this criterion, one effectively excludes disk stars, stars belonging to the Large and Small Magellanic Clouds, the Sagittarius dwarf spheroidal, and other Sagittarius arms. After this matching, our final data set of K-Giant has 6058 sources.
@@ 2.2 RR-Lyrae data set
To build the RR-Lyraes catalogue, we started the process by using the 271779 sources catalogued as RR-Lyraes variables by Gaia. Initially, we performed data cleaning by keeping stars with $ruwe<1.4$, and excluding those lacking metallicity, as well as those with errors in metallicity exceeding the absolute value of the metallicity itself. We then discarded all sources with $E(B-V)>0.3$, and the galactic plane $|b|<10^{\circ}$. These cuts yielded a set of 66610 stars, out of which only 2440 had radial velocity measurements provided by Gaia.
To increase the amount of data with measured radial velocities and metallicities, we performed a series of data cross-matching steps. We use the SEGUE catalogue (Ahn et al., 2012) to complete the radial velocities and metallicities of our dataset using the ones measured by Sloan (cross-match with ID table sdss__\__dr17.x1p5____\_\__ _specobj____\_\__ _gaia__\__dr3____\_\__ _gaia__\__source). In this case, no extra points were incorporated.
Additionally, we utilised the RR-Lyrae catalogue provided by Wang et al. (2022) to complete our catalogue with the sources not present in our data set (288 points where incorporated) or to complete our data with radial velocity and metallicity.
After the cross-matching process, we ended up with a catalogue of 73598 sources, of which 6523 have measured radial velocities.
@@
[…]
To ensure consistency, we followed a similar approach used with the K-Giant. To eliminate known objects from the Harris (1996a) and Drlica-Wagner et al. (2020a) catalogues, we applied an angular mask taking into account the distance to the sources. Additionally, to exclude the Sagittarius stream from our analysis, we employed the cut-off criterion of $|B_{Sgr}|<15^{\circ}$ for $b>0^{\circ}$. Then, we focused our analysis on objects with galactic distances between 30 $30\textrm{ kpc}<R_{gal}<100\textrm{ kpc}$. In the final step, we removed stars that did not satisfy the condition $\mu_{\alpha*}^{2}+(\mu_{\delta}+0.1)^{2}<0.5^{2}$ (Conroy et al., 2021) after performing the solar reflex motion correction to the proper motions. Therefore, the final sample has 2446 sources.
Since our aim is to extract the mass of the dark matter subhalo surrounding the Large and Small Magellanic Cloud, we have performed a transformation of coordinates to a new reference system. This particular coordinate reference system is centred in the centre of mass (CM) of the Magellanic Clouds System (MCS), with the $x$-axis aligned with the direction of the velocity of the centre of mass (see appendix A for details), and it is considered a rest frame. In order to obtain the position and velocity of the centre of mass we have considered that the LMC mass is nine times the SMC mass (Craig et al., 2021).
@@ 3 Theoretical model and Likelihood analysis
We used the theoretical model for stellar wakes from dark matter subhalos proposed by Buschmann et al. (2018), where the authors assumed a Plummer sphere for the density profile of the DM subhalo. The reference system used in this Section corresponds to the one centred in the perturbed (in our case the centre of mass of the Magellanic Clouds System), with the $x$-axis in the direction of the velocity of the dark matter subhalo mass (see Appendix A.1 for details). From the collisionless Boltzmann equation, they derived the time-independent phase-space distribution function in the subhalo rest frame
@@
In the equations, $f_{0}(\bar{v})=\frac{n_{0}e^{-\left(\bar{v}+\bar{v}_{s}\right)^{2}/v_{0}^{2}}}{\pi^{3/2}v_{0}^{3}}$, ${v_{0}=\sqrt{2}\sigma_{v}}$ where $\sigma_{v}$ is the velocity dispersion, $\bar{v}_{s}$ and $M_{s}$ are the DM subhalo velocity and mass respectively, ${R_{s}(M_{s})=1.62\sqrt{M_{s}/(10^{8}\rm{M}_{\odot})}}$ and $n_{0}$ is the star density inside the region of interest. This expression can be easily extended for three different velocity dispersion by including $v_{0x}$ $\left(\sigma_{v_{x}}\right)$, $v_{0y}$ $\left(\sigma_{v_{y}}\right)$ and $v_{0z}$ $\left(\sigma_{v_{z}}\right)$. In this case, the distribution function in the subhalo rest frame can be written as
@@
In order to obtain the mass of the DM subhalo of the MCS, we have performed a statistical analysis using the likelihood function to compare observational data in the new reference system and the theoretical model. This analysis was performed by using only the space data ($3D$) and the phase-space data ($6D$). The un-binned likelihood function for the $6D$ analysis is (Buschmann et al., 2018)
@@
where $N_{d}$ is the number of stars in the region of interest (sphere of radius $R$ centred in the centre of mass of the MCS), $N_{s}$ is the predicted number of stars in the same region and $\theta$ are the fixed parameters of our model, that is, $n_{0}$, $v_{0}$, $R_{s}$ and $\bar{v}_{s}$. For the Plummer sphere and for the distribution function of Eq. (2), $N_{s}(M_{s})$ can be computed as (Buschmann et al., 2018)
@@
To determine the DM subhalo mass we have used a Markov-Chain Monte Carlo (MCMC) method. For this purpose, we utilised the emcee package (Foreman-Mackey et al., 2013). We have considered the function
@@
where $x$ stands for the 3D or 6D analysis. The prior used was $1210<\log_{10}M_{s}<$ 12. The functions $p_{x}$ are presented in Eqs. (4) and (5). We have considered $32$ walkers and check the convergence every $100$ steps. To compute the uncertainties, we have used the percentiles $16^{th}$, $50^{th}$ and $84^{th}$ of the samples in the marginalised distributions (Foreman-Mackey et al., 2013).

### v3

@@ Abstract
Aims. Our goal is to study the gravitational effects caused by the passage of the Large Magellanic Cloud (LMC) in its orbit on the stellar halo of the Milky Way.
Methods. We employed Gaia Data Release 3 to construct a halo tracers dataset consisting of K-giant stars and RR-Lyrae variables. Additionally, we compared the data with a theoretical model to estimate the dark matter subhalo mass.
Results. We have improved the characterisation of the local wake and the collective response due to the LMC’s orbit. We have also estimated for the first time the dark subhalo mass of the LMC to be of the order of $1.7\times 10^{11}$ M, which is comparable to previously reported values in the literature.
@@ 1 Introduction
Dark matter (DM) is central to the standard cosmological model (LCDM), providing gravitational support for the formation of galaxies and systems of galaxies (Mo et al., 2010). Its existence is backed by a plethora of observational data, including galaxies’ rotation curves (Zwicky, 1933; Rubin & Ford, 1970), strong and weak gravitational lensing effects (Massey et al., 2010; Clowe et al., 2006), and even the presence of baryonic acoustic oscillation (Planck Collaboration et al., 2020) in the earlier gravitational wells observed in the cosmic microwave background (CMB).
@@
Despite these successes, we still lack a precise detection either in the laboratory (Bernabei et al., 2022; Aprile et al., 2023; Barberio et al., 2023; Amaré et al., 2022) or by indirect astrophysical observation (Abdalla et al., 2022; Acharyya et al., 2023; Abe et al., 2023). Many theoretical candidates arise out of physics beyond the standard model (BSMP), like weakly interacting massive particles (WIMPS). These massive ($m_{DM}\sim 100$ GeV) particles could weakly interact with nucleons, and therefore their signals have been looked at by several laboratory and accelerator experiments. Also, their annihilation signals could be detectable through $\gamma$-ray emission by high-energy telescopes. Despite a massive experimental effort, DM remains a theoretical hypothesis, albeit one with impressive empirical support.
@@
Other DM candidates could be more massive, such as primordial black holes (PBHs) (Carr & Kühnel, 2020), which were recently constrained with a series of consistency tests. Nowadays, there is still room to be an essential contributor to the DM content but this is limited to small windows in mass (Villanueva-Domingo et al., 2021). Other candidates include massive ultralight particles (ULDMs) that could reach masses as low as $10^{-23}$ eV (Hui et al., 2017). Therefore, the possible DM mass range remains unconstrained today. Additionally, DM particles could interact with themselves, and have or not have spin, and other properties, including their mass, remain elusive.
@@
Depending on the nature of the DM particle, there are relevant changes in the structure and number of DM halos and subhalos (Zavala & Frenk, 2019). For example, some candidates, like warm dark matter (WDM), introduce a cut-off scale in the initial power spectrum of mass fluctuations ($m_{DM}\sim 1$ KeV), and others a scale during the non-linear evolution phase in which the DM particles self-interact (SIDM) (Tulin & Yu, 2018). Both processes change the abundance of DM subhalos and the density profiles of the DM halos in comparison to the predictions of the cold dark matter (CDM) model.
@@
Buschmann et al. (2018) show, using CDM simulations, that the gravitational pull of DM subhalos affects the distribution of stars in galactic halos, and could be used to discover dark subhalos (those without star formation in situ, a precise prediction of the CDM model) and also allow one to test the nature of the DM particles itself. This work simulated a passing DM subhalo’s perturbation to the phase space stellar distribution. Stars are pulled towards the subhalo as it passes, leaving a distinctive feature in the halo stars’ velocity and number density, known as a wake. This phenomenon was previously analytically described by Weinberg (1986) due to the gravitational friction that provokes the orbital decay of the satellite galaxies that inhabit the DM subhalo. There have been some efforts to quantify the magnitude of phase-space perturbations caused by the passage of DM subhalos using simulations and their possible detection (Bazarov et al., 2022).
@@
Garavito-Camargo et al. (2019) used CDM simulations to quantify the impact of the LMC’s passage on the density and kinematics of the Milky Way (MW) DM halo and the observability of these structures in the MW’s stellar halo. Their results indicated a pronounced wake, which could be decomposed in a transient response and a collective one in both the DM and stellar halo distributions. Such an effect was observed for the first time in the MW halo stars by Conroy et al. (2021). The authors studied the effects induced by the Magellanic Clouds System (MCS) merging in selected samples of MW stars with precise Gaia Satellite measurements. This detection, and the increasing availability of stellar data from Gaia DR3 release, has paved the way for a more precise measurement of the effect. Measurements of the wake on the perturber systems of reference will allow pursuit further testing of the DM particle’s nature using this merger data, as was proposed recently by Foote et al. (2023) (in the context of systems of galaxies see Furlanetto & Loeb (2002); Buehler & Desjacques (2023)). Furthermore, Aguilar-Argüello et al. (2022) and Cunningham et al. (2020) conducted studies on the decomposition into spherical harmonics of both density and velocity, respectively, in order to quantify the response of the DM halo to the passage of the Large Magellanic Cloud (LMC).
@@
Based on the findings of Conroy et al. (2021), we have used Gaia Data Release 3 to study the DM subhalo of the Magellanic Clouds. Our code developments will also allow us to apply the methodology to other MW satellite galaxies and globular clusters and even to develop methods of detecting the presence of the dark subhalos predicted by the CDM model.
@@
This work is organised as follows. In Section 2, we present the data samples selection methodology used to identify the effects of the DM subhalo of the LMC on the MW stellar halo. In Section 3, we briefly describe the theoretical model. Meanwhile, we present our results in Section 4. Section 5 presents the conclusions and future perspectives. Finally, we present in Appendix A a description of the coordinates’ transformation, in Appendix B the machine learning algorithm used to estimate radial velocities, and in Appendix C a validation of our method of inferring distances.
@@ 2 Data reduction
We studied the gravitational response of the MW’s halo to the passage of the LMC in its orbit. To achieve this, we used the data from Gaia Data Release 3 (Gaia Collaboration et al., 2023, 2016) and created two catalogues of halo tracers; namely, K-giant stars and RR-Lyrae ones. We followed the steps proposed by Conroy et al. (2021) to address this task.
@@ 2.1 K-giant dataset
To construct the K-giant catalogue, we started the analysis with 162240774 sources characterised by $ruwe$ values below 1.4, parallax measurements lower than 0.2 mas, and galactic latitudes of $|b|>10^{\circ}$ to remove the galactic plane. To ensure data quality, we performed a series of cleaning procedures. First, we eliminated sources lacking proper motion and photometric data. To account for dust extinction, we obtained the dust map from Green (2018) and considered the Schlegel, Finkbeiner &\&& Davis (SFD) map to derive the excess colour, $E(B-V)$. Subsequently, we discarded all sources with $E(B-V)>0.3$. To obtain the corrected magnitudes, we considered the following coefficients: $A_{G}/E(B-V)=2.4$, $A_{BP}/E(B-V)=2.58$, and $A_{RP}/E(B-V)=1.65$. To focus solely on the giant branch, we restricted the selection to sources satisfying the condition $21.4<(BP^{*}-RP^{*})<$ 2, where $BP^{*}$ and $RP^{*}$ represent the corrected magnitudes. Next, following Riello et al. (2021), we performed the $3\sigma$ cut upon the corrected BP and RP flux excess factor ($C^{*}$). After completing the data-cleaning process, to ensure the purity of our catalogue specifically for K-giant stars, we performed a cross-match with the spectral types provided by Gaia (gaiadr3.astrophysical parameters) and obtained a dataset of 490669 sources. Finally, we restricted our analysis to objects within a galactocentric distance between $30$ kpc and $100$ kpc, leaving 245086 sources. Among them, only 10989 had radial velocities measured by Gaia (Katz et al., 2023).
@@
To validate our distance inference method, we compared our calculated distances with reference values for some globular clusters, the Magellanic Clouds (LMC and SMC), and some satellite galaxies (see Appendix C for details). Our method successfully reproduced tabulated distances, since the mean calculated distances differed by less than $20\%$.
Afterwards, we implemented several masks to exclude known objects from our analysis. Specifically, we applied angular and distance masks to all the globular clusters listed in Harris (1996a) and all the satellite galaxies reported in Drlica-Wagner et al. (2020a).
Following the methodology proposed by Conroy et al. (2021), we employed proper motions to eliminate structures linked to the Sagittarius stream. To achieve this, we initially correct the proper motions due to the solar reflex motion, with the Gala package (Price-Whelan, 2017; Price-Whelan et al., 2020). The used parameters were $R_{\odot}=8.122$ kpc (GRAVITY Collaboration et al., 2019), $(V_{R,\odot},V_{\phi,\odot},V_{Z,\odot})=(-12.9,245.6,7.78)$ km/s (Drimmel & Poggio, 2018), and the distance of the Sun from the Galactic mid-plane, $Z_{\odot}=20.8$ pc (Bennett & Bovy, 2019). For $b>10^{\circ}$ and $|B_{Sgr}|<15^{\circ}$, where $B_{Sgr}$ is the latitude in the frame of the Sagittarius orbital plane, we removed part of the northern arm of the stream by taking out the sources that have $\mu_{\alpha*}>-1.3$ mas/yr, $-0.4<\mu_{\delta}<0.3$ mas/yr, and $\mu_{\delta}>1.7\mu_{\alpha*}+0.4$. To eliminate the rest of the north arm, we applied a mask to the region with coordinates $b>0^{\circ}$ and $180^{\circ}<l<210^{\circ}$. The final selection of sources was based on proper motions; that is, we kept only those that satisfied $\mu_{\alpha*}^{2}+(\mu_{\delta}+0.1)^{2}<0.5^{2}$ (Conroy et al., 2021). By implementing this criterion, one effectively excludes disk stars, stars belonging to the Large and Small Magellanic Clouds, the Sagittarius dwarf spheroidal, and other Sagittarius arms. After this matching, our final dataset of K giants had 6058 sources.
@@ 2.2 RR-Lyrae dataset
To build the RR-Lyrae catalogue, we started the process by using the 271779 sources catalogued as RR-Lyrae variables by Gaia. Initially, we performed data cleaning by keeping stars with $ruwe<1.4$, and excluding those lacking metallicity, as well as those with errors in metallicity exceeding the absolute value of the metallicity itself. We then discarded all sources with $E(B-V)>0.3$, and the galactic plane $|b|<10^{\circ}$. These cuts yielded a set of 66610 stars, of which only 2440 had radial velocity measurements provided by Gaia.
To increase the amount of data with measured radial velocities and metallicities, we performed a series of data cross-matching steps. We used the SEGUE catalogue (Ahn et al., 2012) to complete the radial velocities and metallicities of our dataset, using the ones measured by Sloan (cross-match with ID table sdss__\__dr17.x1p5____\_\__ _specobj____\_\__ _gaia__\__dr3____\_\__ _gaia__\__source). In this case, no extra points were incorporated.
@@
To ensure consistency, we followed a similar approach to that used with the K-giants. To eliminate known objects from the Harris (1996a) and Drlica-Wagner et al. (2020a) catalogues, we applied an angular mask, taking into account the distance to the sources. Additionally, to exclude the Sagittarius stream from our analysis, we employed the cut-off criterion of $|B_{Sgr}|<15^{\circ}$ for $b>0^{\circ}$. Then, we focused our analysis on objects with galactic distances between 30 $30\textrm{ kpc}<R_{gal}<100\textrm{ kpc}$. In the final step, we removed stars that did not satisfy the condition $\mu_{\alpha*}^{2}+(\mu_{\delta}+0.1)^{2}<0.5^{2}$ (Conroy et al., 2021) after performing the solar reflex motion correction to the proper motions. Therefore, the final sample has 2446 sources.
Since our aim is to extract the mass of the DM subhalo surrounding the Large and Small Magellanic Clouds, we performed a transformation of coordinates to a new reference system. This particular coordinate reference system is centred on the centre of mass (CM) of the MCS, with the $x$ axis aligned with the direction of the velocity of the CM (see appendix A for details), and it is considered a rest frame. In order to obtain the position and velocity of the CM, we considered the LMC mass to be nine times the SMC mass (Craig et al., 2022).
@@ 3 Theoretical model and likelihood analysis
We used the theoretical model for stellar wakes from DM subhalos proposed by Buschmann et al. (2018), where the authors assumed a Plummer sphere for the density profile of the DM subhalo. The reference system used in this section corresponds to the one centred on the CM of the MCS, with the $x$ axis in the direction of the velocity of the DM subhalo mass (see Appendix A.1 for details). From the collisionless Boltzmann equation, they derived the time-independent phase-space distribution function in the subhalo rest frame
@@
In the equations, $f_{0}(\bar{v})=\frac{n_{0}e^{-\left(\bar{v}+\bar{v}_{s}\right)^{2}/v_{0}^{2}}}{\pi^{3/2}v_{0}^{3}}$, ${v_{0}=\sqrt{2}\sigma_{v}}$ where $\sigma_{v}$ is the velocity dispersion, $\bar{v}_{s}$ and $M_{s}$ are the DM subhalo velocity and mass, respectively, ${R_{s}(M_{s})=1.62\sqrt{M_{s}/(10^{8}\rm{M}_{\odot})}}$, and $n_{0}$ is the star density inside the region of interest. This expression can be easily extended for three different velocity dispersions by including $v_{0x}$ $\left(\sigma_{v_{x}}\right)$, $v_{0y}$ $\left(\sigma_{v_{y}}\right)$, and $v_{0z}$ $\left(\sigma_{v_{z}}\right)$. In this case, the distribution function in the subhalo rest frame can be written as
@@
In order to obtain the mass of the DM subhalo of the MCS, we performed a statistical analysis using the likelihood function to compare observational data in the new reference system and the theoretical model. This analysis was performed by using only the space data ($3D$) and the phase-space data ($6D$). The un-binned likelihood function for the $6D$ analysis is (Buschmann et al., 2018)
@@
where $N_{d}$ is the number of stars in the region of interest (sphere of radius $R$ centred on the CM of the MCS), $N_{s}$ is the predicted number of stars in the same region, and $\theta$ are the fixed parameters of our model; that is, $n_{0}$, $v_{0}$, $R_{s}$, and $\bar{v}_{s}$. For the Plummer sphere and for the distribution function of Eq. (2), $N_{s}(M_{s})$ can be computed as (Buschmann et al., 2018)
@@
To determine the DM subhalo mass, we used a Markov-Chain Monte Carlo (MCMC) method. For this purpose, we utilised the emcee package (Foreman-Mackey et al., 2013). We considered the function
@@
where $x$ stands for the 3D or 6D analysis. The prior used was $1210<\log_{10}M_{s}<$ 12. The functions, $p_{x}$, are presented in Eqs. (4) and (5). We considered $32$ walkers and checked the convergence every $100$ steps. To compute the uncertainties, we used the $16^{th}$, $50^{th}$, and $84^{th}$ percentiles of the samples in the marginalised distributions (Foreman-Mackey et al., 2013).